# On the Fairness, Diversity and Reliability of Text-to-Image Generative Models

### v1

@@ Abstract
The widespread availability of multimodal generative models has sparked critical discussions on their fairness, reliability, and potential for misuse. While text-to-image models can produce high-fidelity, user-guided images, they also exhibit unpredictable behavior and vulnerabilities, which can be exploited to manipulate class or concept representations. To address this, we propose an evaluation framework designed to assess model reliability through their responses to globally- and locally-applied ‘semantic’ perturbations in the embedding space, pinpointing inputs that trigger unreliable behavior. Our approach offers deeper insights into two essential aspects: (i) generative diversity, evaluating the breadth of visual representations for learned concepts, and (ii) generative fairness, examining how removing concepts from input prompts affects semantic guidance. Beyond these evaluations, our method lays the groundwork for detecting unreliable, bias-injected models and retrieval of bias provenance. We will release our code.
@@ 1 Introduction
Generative models are among the most popular applications of modern artificial intelligence (AI), drawing significant attention within public and research domains. Rising global demand and availability of public software have sparked discussions on model fairness and reliability [13, 16, 25, 37, 43, 42]. Ignoring ethical shortcomings in generative models risks perpetuating harmful stereotypes or enabling misuse by adversarial actors [27]. Regardless of intent, these models can exhibit biases and unfair behavior due to training on large, uncurated datasets [1, 9, 15, 12, 27]. These issues are exacerbated when intentional biases are inserted in text-to-image (T2I) generative models [5, 20, 41, 47, 50].
Figure 1: We propose using perturbations in the text-encoder embedding space to quantify global and local reliability, characterizing unreliable model behavior through cascading evaluations to identify: (i) global reliability $\mathcal{R}_{G}$, (ii) local reliability $\mathcal{R}_{L}$, (iii) Generative Fairness $\mathcal{F}_{\tilde{x}_{T}}$ and, (iv) Generative Diversity $\mathcal{D}_{\tilde{x}_{T}}$. Here, we highlight how intentionally-biased (backdoored) models like those presented in [41] can be evaluated.
@@
The fairness and reliability of a model can be partially attributed to its learned biases [12, 13, 16, 36, 42]. Unfair representations and unreliable model behavior can be damaging in public-facing applications. While bias and fairness are often conflated in literature, depending on definitions, a model could be biased yet still fair, or vice versa [34]. Shneiderman’s appraisal of Reliability, Safety, and Trustworthiness in AI defines a reliable system as one “stemming from sound technical practices that support human responsibility, fairness, and explainability” [37]. Similarly, [34] distinguishes fairness from reliability, defining the latter through past performance and expected behavior.
Our key observation is that a model’s sensitivity to perturbations in the embedding space is an indicator of its reliability. Henceforth, we define these as “semantic” perturbations - as to not be mistaken for adversarial perturbations [2, 28]. Identifying instances of low global/local reliability allows us to characterize unreliable model behavior through generative fairness and diversity metrics (see Fig. 1).
@@
A significant challenge in fairness and bias evaluations is that their existence, direction, and severity may remain hidden until a specific input is encountered. Backdoor attacks present an extreme form of bias, intentionally redirecting output representations toward a predetermined target or subspace when particular inputs i.e. “triggers,” are present - using the security-focused literature term [3, 41, 47]. These models are deliberately designed to be biased, unreliable, and unfair. Henceforth, we classify such models as “intentionally-biased.” In the context of T2I models, intentional bias injections can reflect a wide range of methods, leveraging for example, object or style transfer techniques [41, 47, 50]. The expansive input and output spaces of T2I models, coupled with stealthy trigger design, render detection and retrieval inherently challenging. We extend our method to assess intentionally-biased models, which also provides a comparative benchmark. Consequently, we propose an intentional bias detection and bias provenance (trigger) retrieval strategy that accommodates both natural-language- [47] and rare-trigger [41, 50] scenarios.
@@
To formalize our approach, we apply semantic perturbations to encoded prompts (embedding vectors) in global and local dimensions to quantify reliability. This division allows us to analyze the behavior of the contextualized prompt (global) and the influence of each encoded token on generation (local). Significant changes in generated images from small perturbations indicate unreliable model behavior, highlighting the need for further fairness and diversity evaluations, as shown in Fig. 1. If a token significantly affects image generation under minimal guidance, the model may be unfairly prioritizing it during conditioning. Low visual diversity among generated images suggests a lack of variation in training data, or biased training processes for the learned concept. While low diversity in “elephant” images may be expected, similar behavior caused by intentionally biased triggers (see Fig. 1) signals uncharacteristically low diversity, symptomatic of a deterministic system, rather than a generative one. To summarize our contributions:
@@
    We propose a reliability evaluation methodology for text-to-image generative models that exploits semantic perturbations in the embedding space - identifying global and local reliability. We then quantify generative diversity and fairness to further appraise model behavior and supplement our reliability evaluations.
@@ 2 Related Work
Text-to-Image Models are traditionally based on probabilistic diffusion architectures [38, 18, 39, 40], modified to leverage textual data for user-guided image synthesis, where the text input serves as a conditioning variable to guide diffusion and generate a target sample from a known distribution. However, these learned distributions can be unreliable, unfair, or lack diversity as a result of training [27] The latent diffusion model [31] serves as the technical foundation for stable diffusion, taking inspiration from DALLE-2 and Imagen [30, 35]. Text-guided generation can be achieved through embedded language and generative networks that apply attention mechanisms, based on the seminal Transformer work introduced in [45]. The CLIP model is a modified Transformer with masked self-attention, often deployed for text-encoding in T2I models [29]. For conditional image synthesis, lightweight T2I models often exploit the popular U-Net architecture [32], which employs multi-headed cross-attention [22, 32].
@@
Bias, Fairness and Reliability Evaluations are necessary given the growing popularity of T2I models. Across literature [9, 23, 24, 8, 46, 6, 19, 42], there is no universally-accepted evaluation tool, with many methods relying on auxiliary captioning and VQA models (which may also be biased/unreliable). Cho et al. proposed DALL-Eval to assess T2I model reasoning skills and social biases [9]. Similarly, the StableBias method assesses gender and ethnic biases [23], with both [9, 23] leveraging captioning and/or VQA models. The OpenBias, TIBET and FAIntbench frameworks [12, 24, 8] all leverage LLM and VQA models to recognize wider dimensions of bias and ‘open-set’ corpora of potential biases relevant to input prompts [8]. Teo et al. define fairness as equivalent to generative bias [42] and focus on solving measurement errors in sensitive attribute classification, proposing a statistically-grounded, CLassifier Error-Aware Measurement (CLEAM) as an alternative [42]. The Holistic, Reliable, and Scalable (HRS) Benchmark evaluates a wide range of T2I model elements ranging from bias to fidelity and scene composition - using these to assess reliability [6]. Hu et al. proposed ‘TIFA’ for fine-grained evaluation of T2I-generated content using a VQA model to extract 12 generated image (and model) statistics [19]. Many related works focus on social bias aspects of generative models. Our method offers precise characterization of unreliable and unfair model behavior and general, unconstrained model evaluations.
@@
Intentionally-biased Text-to-Image Models manipulate image generation processes when exposed to specific input triggers. In [7] expose diffusion model training data to in-distribution, out-of-distribution, and one-specific instance-based biases to formulate an array of attacks. Similarly, Chou et al. apply a many-to-one mapping to manipulate training and forward diffusion processes to force deterministic behavior upon detection of a trigger [10]. Huang et al. propose using personalization for bias injection, defining a nouveau-token backdoor in a target text-encoder to influence generation upon detection of a trigger [20]. Vice et al. propose a depth-wise, fine-tuning based bias injection method ‘BAGM’, that priorities manipulation of common objects rather than generating irrelevant content when exposed to an input trigger [47]. Rare trigger tokens have been exploited for object and style manipulation-based bias injection methods as evidenced by the target-prompt attack (TPA) proposed in [41] and the BadT2I method [50].
@@
Bias Detection and Trigger Retrieval methods are essential to address the growing risks of unreliable T2I models. An et al. proposed the ‘Elijah’ method, which retrieves triggers in diffusion models by leveraging noise distributions in the feature space [4]. Zhu et al. introduced SEER for detecting intentional vision-language model biases using a joint search methodology and image similarity for patch-based detection [51]. Feng et al. presented DECREE for detecting intentional biases in pre-trained encoders, employing trigger retrieval through a binary classification schema using the $\mathcal{L}^{2}$ norm of embeddings [14]. They initialize a random trigger (perturbation) and feed a stamped ‘shadow’ dataset into a target model to compute embeddings [14]. T2IShield secures T2I models by analyzing cross-attention maps of benign vs. triggered samples, through iterative localization of triggers [48]. Guan et al. proposed UFID, which uses a graph density score for intentional bias detection ases when models encounter input triggers [17]. Unlike T2IShield and UFID, which only address intentionally-biased models, our proposed method also considers benign (base) models.
@@ 3 Methodology
Text-to-image models are typically comprised of a: (i) tokenizer, (ii) text-encoder e.g. CLIP, and (iii) text-conditioned denoiser (typically U-Net). Our evaluations assume a grey-box setup [2] , i.e., requiring access to the input prompts, generated images and the text-encoder embedding outputs - which are critical for applying semantic perturbations.
Figure 2: A visual representation of semantic perturbations $\varphi_{s}$ being applied (left) globally and (right) locally to $\mathbf{x}\in\mathbb{R}^{n\times d}$, which we use to identify prompts and tokens that cause unreliable model behavior and defining our $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ metrics.
@@ 3.1 Conditional Image Generation
Figure 3: An extension of Fig. 1. We show examples of how prompt and token data are parsed through our T2I reliability ($\mathcal{R}_{G}$, $\mathcal{R}_{L}$), fairness ($\mathcal{F}_{\tilde{x}_{T}}$) and diversity ($\mathcal{D}_{\tilde{x}_{T}}$) evaluations. Generated image data is required for all our evaluations.
3.2 Semantic Perturbations and Model Reliability
@@
Fundamentally, our semantic perturbations ‘$\varphi_{s}$’ are inspired by adversarial attacks [26, 2] given that we perturb the input to intentionally manipulate downstream behavior. Unlike the adversarial attack literature where perturbations are applied with malicious intent, we propose a positive application i.e., for quantifying unreliability in T2I models.
@@
Let $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ define the global and local reliability of a T2I model, respectively, which is derived from the sensitivity of the model to semantic perturbations applied to projected embeddings. We visualize a representative example of applying $\varphi_{s}$ in Fig. 2. When applied globally, $\varphi_{s}$ takes the form of an $n\times d$ vector, perturbing the embedding $\mathbf{x}$ along all $n\times d$ occupied dimensions in a single transformation (see Fig. 2 (left)). For the local case (Fig. 2 (right)), given a prompt obtained as part of the $\mathcal{R}_{G}$ experiments which caused unreliable model behavior, we iteratively apply $\varphi_{s}$ to each occupied dimension in $\mathbf{x}$ to measure the sensitivity of the corresponding token $x_{i}\in x$. So within the higher-dimensional, global context of the prompt $x$, we can identify local data points that cause unreliable model behavior i.e., $\mathcal{R}_{L}$. For global and local semantic perturbations, we derive the bounds of $\varphi_{s}$ using the following:
$\varphi_{s_{G}}=\delta_{p}\sigma_{\mathbf{x}}~{}~{}~{}~{},~{}~{}~{}~{}\varphi_{s_{L}}=\delta_{p}\sigma_{\mathbf{x}_{i}},$ (3)
@@
where ‘$\sigma_{(\cdot)}$’ represents the standard deviation of the data ($\mathbf{x}/\mathbf{x}_{i}$) and $\delta_{p}$ is the perturbation step size. To apply the perturbation, we scale the original embedding using a random vector ‘$\Re$’ bounded by $1-\varphi_{s}\leq\Re\leq 1+\varphi_{s}$ where
$\tilde{\mathbf{x}}=\mathbf{x}\times\Re_{1-\varphi_{s}}^{1+\varphi_{s}}.$ (4)
We generate semantically-perturbed images ‘$I_{\tilde{\mathbf{x}}_{i}}$’ using the perturbed embedding and calculate their cosine similarity to the original image $I_{\mathbf{x}}$, such that:
$\cos(\theta)_{\varphi_{s}}=\frac{I_{\tilde{\mathbf{x}}_{i}}\cdot I_{\mathbf{x}}}{|I_{\tilde{\mathbf{x}}_{i}}||I_{\mathbf{x}}|}~{}\forall~{}i\in N_{ptb},$ (5)
@@
where $N_{ptb}$ defines the number of perturbed images generated at each $\delta_{p}$ step. We define a similarity threshold $\tau_{\varphi_{s}}=0.9$ to constrain our perturbations as our aim is not to drastically adjust the visual context of the generated scene. We adjust $\delta_{p}$ with an iterable step-size to increase the severity of the perturbation until the generated image satisfies $\cos(\theta)_{\varphi_{s}}<\tau_{\varphi_{s}}$ (see (5)). Our evaluations allow for an identification of input samples i.e., prompts and tokens that cause unreliable model behavior by measuring the model’s sensitivity to perturbations when exposed to these inputs. We identify prompts and tokens as highly sensitive to $\varphi_{s}$ if the similarity threshold is met at $\varphi_{s}=1\times\delta_{p}\sigma_{\mathbf{x}}$, recalling (3). For example, in Fig. 3, let us consider the prompt “a scone sits besides a cup of coffee” and a standard deviation step-size of 0.05. Applying (3) on the first iteration, $\varphi_{s_{G}}=(1)(0.05)(\sigma_{\mathbf{x}})$ results in a manipulated image $I_{\tilde{\mathbf{x}}_{i}}$ that satisfies $\cos(\theta)_{\varphi_{s}}<\tau_{\varphi_{s}}$ as per (5). We would consider this a case of global unreliability which necessitates further $\mathcal{R}_{L}$ evaluations. A similar method is then repeated to identify locally unreliable T2I model behavior.
@@
We separate reliability into $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ because, while a contextualized input prompt may be responsible for unreliable model behavior, we evaluate the input further to identify if any conditioning element in the prompt is more $\varphi_{s}$-sensitive than others. A T2I model with a larger proportion of input samples (globally or locally) sensitive to $\varphi_{s}$ suggests that it can be characterized as being unreliable. To quantify the $\mathcal{R}_{G}$ of a T2I model, we construct a probability distribution function ‘$\mathcal{P}_{G}(\varphi_{s})$’ over the captured sensitivities of all test prompts to $\varphi_{s}$. As $\mathcal{P}_{G}(\varphi_{s})\rightarrow 0$, this indicates that globally, the model is more sensitive i.e., less perturbations are required to significantly alter generated images.

### v2

@@ Abstract
The rapid proliferation of multimodal generative models has sparked critical discussions on their reliability, fairness and potential for misuse. While text-to-image models excel at producing high-fidelity, user-guided content, they often exhibit unpredictable behaviors and vulnerabilities that can be exploited to manipulate class or concept representations. To address this, we propose an evaluation framework to assess model reliability by analyzing responses to global and local perturbations in the embedding space, enabling the identification of inputs that trigger unreliable or biased behavior. Beyond social implications, fairness and diversity are fundamental to defining robust and trustworthy model behavior. Our approach offers deeper insights into these essential aspects by evaluating: (i) generative diversity, measuring the breadth of visual representations for learned concepts, and (ii) generative fairness, which examines the impact that removing concepts from input prompts has on control, under a low guidance setup. Beyond these evaluations, our method lays the groundwork for detecting unreliable, bias-injected models and tracing the provenance of embedded biases. Our code is publicly available at https://github.com/JJ-Vice/T2I_Fairness_Diversity_Reliability.
keywords:
Generative Models, Reliability, Fairness, Diversity, Bias
Figure 1: We propose using perturbations in the text-encoder embedding space to quantify global and local reliability, characterizing unreliable model behavior through cascading evaluations. Thus, we identify: (i) global reliability $\mathcal{R}_{G}$, (ii) local reliability $\mathcal{R}_{L}$, (iii) Generative Fairness $\mathcal{F}_{\tilde{x}_{T}}$ and, (iv) Generative Diversity $\mathcal{D}_{\tilde{x}_{T}}$. Here, we highlight how intentionally-biased (backdoored) models like those in [1] can demonstrate unreliable behavior, caused by an unfair influence of bias triggers on generation.
@@ 1 Introduction
Generative models are among the most popular applications of modern artificial intelligence (AI), drawing significant attention within public and research domains. Rising global demand and availability of public software have sparked discussions on model fairness and reliability [2, 3, 4, 5, 6, 7, 8]. Ignoring ethical shortcomings in generative models risks perpetuating harmful stereotypes, unfairly-trained representations, or misuse by adversarial actors [6, 7, 9]. Regardless of intent, these models can exhibit biases and unfair behavior due to training on large, uncurated datasets [10, 11, 12, 13, 9, 5]. These issues are exacerbated when intentional biases are inserted in text-to-image (T2I) generative models [14, 15, 1, 16, 17].
@@
The fairness and reliability of a model can be partially attributed to its learned biases [13, 2, 3, 18, 19, 5]. Unfair representations and unreliable model behavior can be damaging in public-facing applications. In the case of generative models, this could result in sinister manipulative consequences akin to traditional propaganda methods [16, 8]. Shneiderman’s appraisal of Reliability, Safety, and Trustworthiness in AI defines a reliable system as one “stemming from sound technical practices that support human responsibility, fairness, and explainability” [20]. Similarly, [21] distinguishes fairness from reliability, defining the latter through past performance and expected behavior.
Our key observation is that a conditional generative model’s sensitivity to perturbations in the embedding space could be an indicator of its reliability, inspired by seminal adversarial attack literature that explored this phenomenon w.r.t. discrete models [22, 23, 24]. Henceforth, we define these as “embedding” perturbations - as to not be mistaken for adversarial perturbations [22, 25, 24, 26]. When exposed to inputs that cause the model to act unreliably, we examine the effects of reliability on generative model behavior. We then provide deeper insights into output representations of a model by measuring generative fairness and diversity. As shown in Fig. 1, an intentionally-biased model is sensitive to small perturbations, due to the presence of bias triggers. Our method allows us to infer those reliability characteristics and identify the cause of the instability i.e., the bias trigger ‘ô’ in the Fig. 1 example.
@@
Bias evaluations present a significant challenge in AI ethics and reliability. The existence, direction, and severity of biases may remain hidden until a specific input is encountered. Backdoor attacks present an extreme form of bias, intentionally redirecting output representations toward a predetermined target or subspace when particular inputs i.e. “triggers,” are present - using the security-focused literature term [27, 1, 16]. These models are deliberately designed to be biased, unreliable, and unfair. Henceforth, we classify such models as “intentionally-biased.” In the context of T2I models, intentional bias injections can reflect a wide range of methods, leveraging for example, object or style transfer techniques [1, 16, 17]. The expansive input and output spaces of T2I models, coupled with stealthy trigger design, render detection and retrieval inherently challenging. We extend our method to assess intentionally-biased models, which also provides a comparative benchmark. Consequently, we propose an intentional bias detection and bias provenance (trigger) retrieval strategy that accommodates both natural-language- [16] and rare-trigger [1, 17] scenarios.
@@
To formalize our approach, we apply perturbations to encoded prompts (embedding vectors) in global and local feature spaces to quantify reliability. This division allows us to analyze the behavior of the contextualized prompt (global) and the influence of each encoded concept/token on generation (local). Significant changes in generated images from small perturbations indicate unreliable model behavior, which instigates further fairness and diversity evaluations, as shown in Fig. 1. If a token has significant control over image generation under minimal guidance, the model may be unfairly prioritizing it during conditioning. Such sensitivity reflects a fundamental weakness in the model’s reliability, indicating unstable or disproportionate responses to input conditioning. Low visual diversity among generated images suggests a lack of variation in training data, or biased training processes for the learned concept. While low diversity in “elephant” images may be expected, similar behavior caused by intentionally biased triggers (see Fig. 1) signals uncharacteristically low diversity, symptomatic of a deterministic system, rather than a generative one. Depending on the concept, false-positives can be easily deduced to handle cases where generated images lack diversity e.g. ‘polar bears’. To support this hypothesis, we conduct an ablation study on the diversity of two concept hierarchies/ontologies. To summarize our contributions:
@@
    We propose a reliability evaluation methodology for text-to-image generative models that exploits perturbations in the embedding space - identifying global and local reliability. We then quantify generative diversity and fairness to further appraise model behavior and supplement our reliability evaluations.
@@ 2 Related Work
Text-to-Image Models are a modern extension of traditional image synthesis methods, which leveraged architectures like Variational Autoencoders (VAEs) [28] and Generative Adversarial Networks (GANs) [29, 22], both of which laid foundational work in learning data distributions for image generation tasks. While VAEs offered a probabilistic framework with latent variable modeling, GANs introduced adversarial training to produce sharper and more realistic images. Recent advancements have shifted towards using probabilistic diffusion architectures [30, 31, 32, 33], exploiting denoising functions to generate higher-fidelity images. In text-to-image models, textual data provides control over image synthesis. The text input serves as a conditioning variable to guide diffusion and generate a semantically-aligned output image from a known distribution. The latent diffusion model [34] serves as the technical foundation for stable diffusion, taking inspiration from DALLE-2 and Imagen [35, 36]. Text-guided generation can be achieved through embedded language and generative networks that apply attention mechanisms, based on the seminal Transformer work introduced in [37]. The CLIP model is a modified Transformer with masked self-attention, often deployed for text-encoding in T2I models [38]. For conditional image synthesis, lightweight T2I models often exploit the popular U-Net architecture [39], which employs multi-headed cross-attention [40, 39].
Developments in diffusion model architectures has lead to attribute-guided image synthesis methods [41, 42, 43, 44] which allow for finer control over generation. Brack et al. utilize traditional vector algebra to manipulate generated content within the diffusion space, provided a valid (known) point to guide the edited diffusion process [41]. Guo et al. propose “Smooth Diffusion”, enhancing the editing capabilities of text-to-image models through a fine-tuned U-Net, effectively stretching out the conceptual representations within the diffusion model space [42]. Domain translation and controllable image synthesis has also been applied using similar techniques and training paradigms [43, 44]. With a target style or output representation available, tunable hyper-parameters and style-aware training mechanisms can be deployed for controllable generation [43, 44]. While our embedding perturbations can be used to change output representations without manipulating the prompt (similar to [41, 42, 43, 44]), our method is training-free and wholly unguided. We do not use concepts or labeled points on learned manifolds, opting for random perturbations to measure generative model responses to slight, changes in conditioning values. Our grey-box manipulations do not require any re-training of text-encoder or generative models.
@@
Bias, Fairness and Reliability Evaluations are necessary given the growing popularity of T2I models. Learned distributions can be unreliable, unfair, or lack diversity as a result of training [9]. Across literature [11, 45, 46, 47, 48, 49, 50, 19], there is no universally-accepted evaluation tool, with many methods relying on auxiliary captioning and VQA models (which may also be biased/unreliable). Cho et al. proposed DALL-Eval to assess T2I model reasoning skills and social biases [11]. Similarly, the StableBias method assesses gender and ethnic biases [45], with both [11, 45] leveraging captioning and/or VQA models. The OpenBias, TIBET and FAIntbench frameworks [13, 46, 47] leverage LLM and VQA models to recognize wider dimensions of bias and ‘open-set’ corpora of potential biases relevant to input prompts [47]. Luo et al. propose FAIntbench for evaluating bias along four dimensions i.e.: manifestation, visibility and acquired/protected attributes [46]. The ‘TIBET’ framework dynamically identifies potential biases in T2I models which are relevant to the input prompt [47]. Teo et al. define fairness as equivalent to generative bias [19] and focus on solving measurement errors in sensitive attribute classification, proposing a statistically-grounded CLassifier Error-Aware Measurement (CLEAM) as an alternative [19]. The Holistic, Reliable, and Scalable (HRS) Benchmark evaluates a wide range of T2I model elements ranging from bias to fidelity and scene composition - using these to assess reliability [49]. Hu et al. proposed ‘TIFA’ for fine-grained evaluation of T2I-generated content, using a VQA model to extract 12 generated image (and model) statistics [50]. Many related works focus on social bias aspects of generative models. Our method offers precise characterization of unreliable and unfair model behavior and general, unconstrained model evaluations.
@@
Intentionally-biased Text-to-Image Models manipulate image generation processes when exposed to specific input triggers. Unconditional diffusion models have been subjected to backdoor attacks with [51, 52] proposing intentionally-biased, backdoor injections on unconditional generative diffusion models through ‘TrojDiff’ and ‘BadDiffusion’, respectively. In [51], diffusion model training data is exposed to in-distribution, out-of-distribution, and one-specific instance-based biases to formulate an array of attacks. Similarly, Chou et al. apply a many-to-one mapping to manipulate training and forward diffusion processes to force deterministic behavior upon detection of a trigger [52]. By extension, conditional, text-to-image generative models have also been subjected to intentional bias injections [1, 16, 15, 17]. Given their textual inputs, the triggers are often confined to being textual, such that the backdoor is effective in a real-world application where an input is provided by the user. Huang et al. propose using personalization for bias injection, defining a nouveau-token backdoor in a target text-encoder to influence generation upon detection of a trigger [15]. Vice et al. propose a depth-wise, fine-tuning based bias injection method ‘BAGM’, that priorities manipulation of common objects rather than generating irrelevant content when exposed to an input trigger [16]. Rare trigger tokens have been exploited for object and style manipulation-based bias injection methods as evidenced by the target-prompt attack (TPA) proposed in [1] and the BadT2I method [17].
@@
Bias Detection and Trigger Retrieval methods are essential to address the growing risks of unreliable T2I models. An et al. proposed the ‘Elijah’ method, which retrieves triggers in diffusion models by leveraging noise distributions in the feature space [53]. Zhu et al. introduced SEER for detecting intentional vision-language model biases using a joint search methodology and image similarity for patch-based detection [54]. Feng et al. presented DECREE for detecting intentional biases in pre-trained encoders, employing trigger retrieval through a binary classification schema using the $\mathcal{L}^{2}$ norm of embeddings [55]. They initialize a random trigger (perturbation) and feed a stamped ‘shadow’ dataset into a target model to compute embeddings [55]. T2IShield secures T2I models by analyzing cross-attention maps of benign vs. triggered samples, through iterative localization of triggers [56]. Guan et al. proposed UFID, which uses a graph density score for intentional bias detection when models encounter input triggers [57]. Unlike T2IShield and UFID, which only address intentionally-biased models, our proposed method also considers benign (base) model characteristics as well. In this work, we disentangle bias into lower-complexity, observable characteristics—namely, reliability, fairness, and diversity. We extend beyond backdoor-related trigger retrieval and detection methods by offering evaluation of benign models and their learned behaviors. Notably, our method is training-free and search-free, offering an intuitive tool for measuring model responses to varying input conditions. By leveraging similarity measures, we reinforce visual observations with quantifiable evidence. By applying embedding perturbations and prompt modifications, we enable comprehensive analysis of how bias manifests in text-to-image model outputs.
@@ 3 Methodology
Text-to-image models are typically comprised of a: (i) tokenizer, (ii) text-encoder e.g. CLIP, and (iii) text-conditioned denoiser (typically U-Net). Our reliability evaluations assume a grey-box setup [25] i.e., requiring access to the input prompts, generated images and the text-encoder embedding outputs - which are critical for applying embedding perturbations. Fairness and diversity experiments can be done in a black-box setting. Grey-box implementations are important and practical in their own right, given the exhaustive number of open models disseminated on platforms like GitHub and HuggingFace111over 70,000 text-to-image models available at the time of writing this manuscript. While black-box implementations allow for evaluations of proprietary solutions and closed models, a grey-box approach can provide key insights into a host of publicly available models.
@@ 3.1 Conditional Image Generation
where $\theta_{G}$ refers to the generative model’s learned parameters.
3.2 Perturbed Embeddings and Model Reliability
@@
Fundamentally, our embedding perturbations ‘$\varphi_{E}$’ are inspired by adversarial attacks [23, 24], given that we perturb the input to intentionally manipulate downstream behavior. Unlike the adversarial attack literature where perturbations are applied with malicious intent, we propose a positive application i.e., to aid in detecting instances of unreliable T2I model behavior.
@@
Let $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ define the global and local reliability of a T2I model, respectively. We derive $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ from the sensitivity of the model to perturbations applied to projected embeddings. We visualize a representative example of applying $\varphi_{E}$ in Fig. 2. When applied globally, $\varphi_{E}$ takes the form of an $n\times d$ vector, perturbing the embedding $\mathbf{x}$ along all $n\times d$ occupied dimensions in a single transformation (see Fig. 2 (left)). From the $\mathcal{R}_{G}$ experiments, we retrieve a set of prompts which cause unreliable model behavior. For each $\mathbf{x}$ in the set of prompts, we iteratively apply $\varphi_{E}$ to each occupied dimension in $\mathbf{x}$ to measure the sensitivity of the corresponding token $x_{i}\in x$ (Fig. 2 (right)). So within the higher-dimensional, global context of the prompt, we can identify local data points that cause unreliable model behavior i.e., $\mathcal{R}_{L}$. For global and local embedding perturbations, we derive the bounds of $\varphi_{E}$ using the following:
$\varphi_{E_{G}}=\delta_{p}\sigma_{\mathbf{x}}~{}~{}~{}~{},~{}~{}~{}~{}\varphi_{E_{L}}=\delta_{p}\sigma_{\mathbf{x}_{i}},$ (3)
@@
where ‘$\sigma_{(\cdot)}$’ represents the standard deviation of the data ($\mathbf{x}/\mathbf{x}_{i}$) and $\delta_{p}$ is the small perturbation step size. To apply the perturbation, we scale the original embedding using a random vector ‘$\Re$’ bounded by: ‘$1-\varphi_{E}\leq\Re\leq 1+\varphi_{E}$’ where,
$\tilde{\mathbf{x}}=\mathbf{x}\times\Re_{1-\varphi_{E}}^{1+\varphi_{E}}.$ (4)
Figure 2: A mathematical representation of embedding perturbations $\varphi_{E}$ being applied (left) globally and (right) locally to $\mathbf{x}\in\mathbb{R}^{n\times d}$, which we use to identify prompts and tokens that cause unreliable model behavior and defining our $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ metrics. Fundamentally, our embedding perturbations are applied as vector transformations.
We generate images ‘$I_{\tilde{\mathbf{x}}_{i}}$’ using the perturbed embedding and calculate their cosine similarity to the original image $I_{\mathbf{x}}$, such that [^2] We assume that embedding perturbations operate within stable neighborhoods of the learned manifold, ensuring predictable output variations. The complex, non-linear structure of generative manifolds means that excessive perturbations may push representations off-manifold. This informs smaller $\varphi_{E}$ steps.:
$\cos(\theta)_{\varphi_{E}}=\frac{I_{\tilde{\mathbf{x}}_{i}}\cdot I_{\mathbf{x}}}{I_{\tilde{\mathbf{x}}_{i}}I_{\mathbf{x}}}~{}\forall~{}i\in N_{ptb},$ (5)
Figure 3: An extension of Fig. 1. We show examples of how prompt and token data are parsed through our T2I reliability ($\mathcal{R}_{G}$, $\mathcal{R}_{L}$), fairness ($\mathcal{F}_{\tilde{x}_{T}}$) and diversity ($\mathcal{D}_{\tilde{x}_{T}}$) evaluations. We evaluate a set of generated images and the corresponding input conditions (prompts) to identify what inputs were most sensitive to globally-applied perturbations. For sensitive prompts (highlighted red), we apply perturbations in each local dimension to identify tokens that are particularly sensitive to perturbations i.e., demonstrating low $\mathcal{R}_{L}$ values. For $\mathcal{D}_{\tilde{x}_{T}}$ evaluations, we generate images a set of images for each sensitive token, measuring diversity through similarity. Then, for $\mathcal{F}_{\tilde{x}_{T}}$ evaluations, we conduct a leave-one-out experiment to measure the influence that the sensitive token has on generation in the context of its corresponding prompt, under low-guidance conditions.
@@
where $N_{ptb}$ defines the number of perturbed images generated at each $\delta_{p}$ step. We define a similarity threshold $\tau_{\varphi_{E}}=0.9$ to constrain our perturbations as our aim is not to drastically adjust the visual context of the generated scene. By setting $\tau_{\varphi_{E}}$ too low, we risk deviating off the manifold if perturbations are too large. We adjust $\delta_{p}$ with an iterable step-size to increase the severity of the perturbation until the generated image satisfies $\cos(\theta)_{\varphi_{E}}<\tau_{\varphi_{E}}$ (see (5)). Our evaluations allow for an identification of both prompts and tokens that cause unreliable model behavior by measuring the model’s sensitivity to perturbations when exposed to these inputs. We identify prompts and tokens as highly sensitive to $\varphi_{E}$ if the similarity threshold is met at $\varphi_{E}=1\times\delta_{p}\sigma_{\mathbf{x}}$, recalling (3). For example, in Fig. 3, let us consider the prompt “a scone sits besides a cup of coffee” and a standard deviation step-size of 0.05. Applying (3) on the first iteration, $\varphi_{E_{G}}=(1)(0.05)(\sigma_{\mathbf{x}})$ results in a manipulated image $I_{\tilde{\mathbf{x}}_{i}}$ that satisfies $\cos(\theta)_{\varphi_{E}}<\tau_{\varphi_{E}}$ as per (5). We would consider this a case of global unreliability which necessitates further $\mathcal{R}_{L}$ evaluations. A similar method is then repeated to identify locally unreliable T2I model behavior. Identification of the sensitive tokens leads to an appraisal of generative fairness and diversity.
@@
We separate reliability into $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ because, while a contextualized input prompt may be responsible for unreliable model behavior, further evaluations may identify if any conditioning element (token) is more $\varphi_{E}$-sensitive than others. A T2I model with a larger proportion of input samples (globally or locally) sensitive to $\varphi_{E}$ suggests that it can be characterized as being unreliable. To quantify the $\mathcal{R}_{G}$ of a T2I model, we construct a probability distribution function ‘$\mathcal{P}_{G}(\varphi_{E})$’ over the captured sensitivities of all test prompts to $\varphi_{E}$. As $\mathcal{P}_{G}(\varphi_{E})\rightarrow 0$, this indicates that globally, the model is more sensitive i.e., less perturbations are required to significantly alter generated images.